# Opportunities and Challenges of Large Language Models for Low-Resource Languages in Humanities Research

### v3

@@ 7.6.1 Gaps Between Available Corpora, Model Capabilities, and Research Needs
Availability of Corpora The available corpora for low-resource languages are diverse, but they are also fraught with issues. Sources include digitization projects of historical documents, language archives from specific regions, and limited academic research collections. These sources remain scattered and significantly smaller in scale compared to high-resource languages. While expanded context windows in models like GPT-5 (400K tokens) and Claude Sonnet 4 (1M tokens) can accommodate entire small language corpora [6], this does not solve the fundamental problem of data quality. The lack of standardized curation processes leads to numerous annotation errors and irregular records, which even sophisticated architectures cannot fully compensate for. Furthermore, many low-resource languages exist primarily in oral traditions, and their digital presence remains minimal despite recent digitization efforts.
Model Capabilities Large language models have evolved significantly in addressing low-resource language challenges, yet fundamental limitations persist. While GPT-5’s unified reasoning architecture can dynamically adjust computational depth based on linguistic complexity [105], and Claude 4’s extended thinking enables thorough analysis of limited linguistic data [7], these advances primarily benefit languages with at least moderate digital presence.
The introduction of MoE architectures allows for language-specific parameter allocation [157], preventing interference between high and low-resource languages. However, the routing mechanisms themselves are trained predominantly on high-resource language patterns, potentially misallocating computational resources for truly low-resource languages. Chain-of-thought reasoning has shown impressive results—achieving 33.8% performance gains with just 1,000 reasoning traces in languages like Swahili [144]—but this assumes sufficient conceptual overlap between pivot and target languages, an assumption that fails for linguistically isolated languages.
Safe-completion frameworks in GPT-5 help prevent hallucination when operating at the edge of model knowledge [104], yet this often results in models being overly cautious with low-resource languages, refusing to generate potentially useful outputs due to uncertainty. The models still struggle with unique grammatical structures, tonal systems, and non-concatenative morphology common in many low-resource languages.
@@ 8.1 Large Language Models in Linguistic Research: the Pilot and the Copilot
Concrete examples make this distinction clearer. The participatory machine translation initiative across Africa, known as Masakhane [106], illustrates how collective engagement of local experts in sampling, labeling, and evaluation produces improvements on several dimensions simultaneously: higher annotation quality, greater reusability of data across projects, and increased social legitimacy of downstream applications. This is not simply a matter of accuracy. The collaborative loop fosters community trust and ensures that the resulting resources align with local priorities and linguistic realities—an essential criterion in domains where language is bound up with identity and sovereignty. A similar logic underpins documentation workflows such as GlossLM [43], which leverages cross-lingual pretraining on interlinear glossed text to provide word-by-word suggestions across thousands of under-documented languages. The system may propose morpheme boundaries, glosses, or candidate syntactic structures, thereby accelerating tedious steps. Yet, the final quality is not determined by the system itself: it depends on human adjudication and on the continual refinement of annotation codebooks. In this sense, the workflow forms a closed feedback loop: copilot suggestion, expert adjudication, codebook feedback, improved copilot suggestions. What emerges is not a substitution of human expertise but a cyclical co-evolution of human conventions and machine affordances.
@@
Then, when might an LLM legitimately assume the role of pilot? Importantly, this decision cannot rest on intuitive judgments about model fluency or perceived intelligence. Instead, it requires falsifiable criteria. The “Alternative Annotator Test” (AAT) [17] offers such a criterion: a model may replace or precede human annotators only when (i) pre-established quality and cost thresholds are met within a well-defined domain, and (ii) the model’s failure modes are both interpretable and constrained. This test treats the model as one more annotator whose reliability must be benchmarked against inter-annotator agreement among humans [9, 14]. A model that passes AAT is not assumed to be perfect, but to function as an alternative annotator whose strengths and weaknesses are transparent, auditable, and integrated into governance protocols.
@@
This same rationale applies beyond annotation and into analytic domains such as diachronic semantics [50, 124]. Consider workflows that combine LLM-driven word-sense induction (WSI) over temporal slices with regression analysis of sense proportions as functions of time, speaker age, and institutional role [4, 44, 18]. The copilot mode allows models to generate candidate sense clusters, align them across historical periods, and surface statistical discrepancies, while human reviewers validate or challenge these outputs. Such a division of labor makes possible new scales of inquiry while preserving scholarly accountability. The point is not to outsource interpretation but to reconfigure it: the model becomes a hypothesis generator, while the human expert adjudicates theoretical import. Evidence from large-scale studies reinforces this point. A recent paper in PNAS [63] analyzing 7.9 million U.S. Congressional speeches (1873–2010) demonstrated that semantic change is not primarily generational. Instead, speakers across all age groups adapt to contemporary semantic trends, with older speakers typically lagging behind by only two to three years and, in some cases, even leading sense innovation [64]. Such findings underscore the necessity of rigorous, large-scale methods and also highlight why interpretability and transparency in LLM-driven pipelines are indispensable: without careful human-guided validation, one risks mistaking spurious statistical shifts for genuine semantic change.
@@ 8.2.1 Automatic recognition and modeling of language variants and dialect differences
GeoAdapt, a recent approach [55], demonstrates the effectiveness of injecting geolinguistic signals directly into encoder architectures. Instead of relying on a single global classifier, which risks overgeneralization and false positives, GeoAdapt leverages region-conditioned plausible language sets [137]. This strategy narrows the hypothesis space to dialects actually spoken in a given region, significantly reducing misclassification rates at scale. For example, in multilingual contexts such as Switzerland, studies on Swiss German reveal that while there is a robust and genuine dialectal signal detectable in embeddings, the proximity of closely related varieties leads to persistent confusion [132]. This highlights an important methodological pitfall: reporting only macro-averaged accuracy across languages obscures dialect-level failures and paints an overly optimistic picture of performance.
Dialect-sensitive modeling is equally crucial in settings involving code-switching. While LLMs now consistently outperform strong supervised baselines on benchmarked code-switching tasks [66], their performance deteriorates when confronted with more fluid, naturally occurring mixing patterns or when faced with domain shifts such as informal social media text. This brittleness underscores the need for stress-testing models under realistic multilingual conditions, rather than relying solely on curated datasets.
In the Arabic NLP community, the introduction of AraDiCE provides a concrete response to these challenges. Unlike many prior resources that implicitly privilege Modern Standard Arabic (MSA), AraDiCE explicitly pairs dialectal coverage with cultural knowledge [100]. This design avoids “standard-language centrism,” ensuring that dialects are treated as primary linguistic objects rather than deviations. Moreover, AraDiCE introduces quantifiable measures of coverage across dialect regions, allowing researchers to more transparently assess which varieties are well supported and which remain underrepresented [100].
@@
  • Audit and control for toponym shortcuts: models may learn to rely on explicit place names (“Cairo,” “Zurich”) rather than genuine linguistic cues. To mitigate this, it is preferable to use coarse regional metadata, and when necessary, remove place names altogether during training and evaluation [55].
  • Incorporate code-switch stress testing by default: model evaluation should include tasks with rapid and fluid switching, reflecting authentic multilingual discourse rather than only balanced, neatly segmented datasets [66].
  • Stratify metrics within a language by variety or region: rather than reporting only cross-language means, which may obscure intra-language disparities, fine-grained stratification reveals performance trade-offs and highlights which dialects remain underserved [38].
@@ 8.2.2 Internal mechanisms and cross-language generalization: a multilingual core and monolingual regions
Multilingual Transformers exhibit a sparse, modular organization: a compact multilingual core carries language-agnostic structure and meaning, while language-/family-specific regions capture orthography, morphology, and idiosyncratic constructions [41, 111, 128]. This motivates freeze-and-specialize: keep the core stable and attach small per-language adapters/LoRA or sparsely routed experts to absorb variation, reducing cross-language interference under fixed compute and mitigating the “curse of multilinguality” [110, 126, 39, 8].
@@
  • Representations: cross-script pairs align in core layers; clustering follows syntax/semantics rather than script [111, 20].
  • Causality: disabling locals hurts orthography/morphology with minor effect on cross-lingual semantics; core perturbations show the converse [110, 41].
  • Routing/sparsity: MoE gates differentiate languages; localized circuits emerge for language-specific phenomena [126, 39].
  • Dynamics: full finetuning forgets more; frozen cores + small locals yield a better stability–specialization trade-off [139].
@@
    Localize: fit small per-language & family modules; for MoE use top-$k$ routing and gate-entropy penalties; for adapters & LoRA enforce sparsity & low rank [126, 39, 110].
@@
    Consolidate (optional): distill core and locals into a sparse expert layer to retain modularity at low latency [148].
@@
Script & morphology engineering often dominates scale: (i) reversible, traceable normalization & transliteration with provenance [140]; (ii) deliberate tokenization (byte & char & subword) by script and typology [145]; (iii) segmentation & reinflection as process-level checks so gains reflect structure, not orthography [126, 39, 139].
@@ 9.1 Development of Specialized Language Models for Low-Resource Languages
    Few-shot and zero-shot learning provide avenues for adapting models to low-resource languages with minimal labeled data. Beyond common approaches like prompt-based learning, innovative frameworks such as Model-Agnostic Meta-Learning (MAML) [40] optimize models for task adaptability, enabling them to perform well with only a handful of examples. Dynamic task re-weighting can prioritize features unique to low-resource languages, ensuring better alignment with linguistic structures. Additionally, task-conditioning embeddings that incorporate linguistic typology can guide models to focus on crucial language-specific attributes. For zero-shot learning, retrieval-augmented reasoning can dynamically fetch relevant multilingual examples, improving the model’s context-aware adaptation to unseen languages.

### v4

@@ 7.6.1 Gaps Between Available Corpora, Model Capabilities, and Research Needs
Availability of Corpora The available corpora for low-resource languages are diverse, but they are also fraught with issues. Sources include digitization projects of historical documents, language archives from specific regions, and limited academic research collections. These sources remain scattered and significantly smaller in scale compared to high-resource languages. While expanded context windows in models like GPT-5 (400K tokens) and Claude Sonnet 4 (1M tokens) can accommodate entire small language corpora [anthropic2025context], this does not solve the fundamental problem of data quality. The lack of standardized curation processes leads to numerous annotation errors and irregular records, which even sophisticated architectures cannot fully compensate for. Furthermore, many low-resource languages exist primarily in oral traditions, and their digital presence remains minimal despite recent digitization efforts.
Model Capabilities Large language models have evolved significantly in addressing low-resource language challenges, yet fundamental limitations persist. While GPT-5’s unified reasoning architecture can dynamically adjust computational depth based on linguistic complexity [openai2025gpt5], and Claude 4’s extended thinking enables thorough analysis of limited linguistic data [anthropic2025claude4], these advances primarily benefit languages with at least moderate digital presence.
The introduction of MoE architectures allows for language-specific parameter allocation [zhou2022moe], preventing interference between high and low-resource languages. However, the routing mechanisms themselves are trained predominantly on high-resource language patterns, potentially misallocating computational resources for truly low-resource languages. Chain-of-thought reasoning has shown impressive results—achieving 33.8% performance gains with just 1,000 reasoning traces in languages like Swahili [yeo2025longcot]—but this assumes sufficient conceptual overlap between pivot and target languages, an assumption that fails for linguistically isolated languages.
Safe-completion frameworks in GPT-5 help prevent hallucination when operating at the edge of model knowledge [openai2025system], yet this often results in models being overly cautious with low-resource languages, refusing to generate potentially useful outputs due to uncertainty. The models still struggle with unique grammatical structures, tonal systems, and non-concatenative morphology common in many low-resource languages.
@@ 8.1 Large Language Models in Linguistic Research: the Pilot and the Copilot
Concrete examples make this distinction clearer. The participatory machine translation initiative across Africa, known as Masakhane [orife2020masakhanemachinetranslation], illustrates how collective engagement of local experts in sampling, labeling, and evaluation produces improvements on several dimensions simultaneously: higher annotation quality, greater reusability of data across projects, and increased social legitimacy of downstream applications. This is not simply a matter of accuracy. The collaborative loop fosters community trust and ensures that the resulting resources align with local priorities and linguistic realities—an essential criterion in domains where language is bound up with identity and sovereignty. A similar logic underpins documentation workflows such as GlossLM [ginn2024glosslmmassivelymultilingualcorpus], which leverages cross-lingual pretraining on interlinear glossed text to provide word-by-word suggestions across thousands of under-documented languages. The system may propose morpheme boundaries, glosses, or candidate syntactic structures, thereby accelerating tedious steps. Yet, the final quality is not determined by the system itself: it depends on human adjudication and on the continual refinement of annotation codebooks. In this sense, the workflow forms a closed feedback loop: copilot suggestion, expert adjudication, codebook feedback, improved copilot suggestions. What emerges is not a substitution of human expertise but a cyclical co-evolution of human conventions and machine affordances.
@@
Then, when might an LLM legitimately assume the role of pilot? Importantly, this decision cannot rest on intuitive judgments about model fluency or perceived intelligence. Instead, it requires falsifiable criteria. The “Alternative Annotator Test” (AAT) [calderon-etal-2025-alternative] offers such a criterion: a model may replace or precede human annotators only when (i) pre-established quality and cost thresholds are met within a well-defined domain, and (ii) the model’s failure modes are both interpretable and constrained. This test treats the model as one more annotator whose reliability must be benchmarked against inter-annotator agreement among humans [artstein-poesio-2008-survey, Braylan_2022]. A model that passes AAT is not assumed to be perfect, but to function as an alternative annotator whose strengths and weaknesses are transparent, auditable, and integrated into governance protocols.
@@
This same rationale applies beyond annotation and into analytic domains such as diachronic semantics [hamilton-etal-2016-diachronic, schlechtweg-etal-2020-semeval]. Consider workflows that combine LLM-driven word-sense induction (WSI) over temporal slices with regression analysis of sense proportions as functions of time, speaker age, and institutional role [amrami2019bettersubstitutionbasedwordsense, Giulianelli_2020, dicarlo2019trainingtemporalwordembeddings]. The copilot mode allows models to generate candidate sense clusters, align them across historical periods, and surface statistical discrepancies, while human reviewers validate or challenge these outputs. Such a division of labor makes possible new scales of inquiry while preserving scholarly accountability. The point is not to outsource interpretation but to reconfigure it: the model becomes a hypothesis generator, while the human expert adjudicates theoretical import. Evidence from large-scale studies reinforces this point. A recent paper in PNAS [doi:10.1073/pnas.2426815122] analyzing 7.9 million U.S. Congressional speeches (1873–2010) demonstrated that semantic change is not primarily generational. Instead, speakers across all age groups adapt to contemporary semantic trends, with older speakers typically lagging behind by only two to three years and, in some cases, even leading sense innovation [kamath.2025.semantic]. Such findings underscore the necessity of rigorous, large-scale methods and also highlight why interpretability and transparency in LLM-driven pipelines are indispensable: without careful human-guided validation, one risks mistaking spurious statistical shifts for genuine semantic change.
@@ 8.2.1 Automatic recognition and modeling of language variants and dialect differences
GeoAdapt, a recent approach [hofmann-etal-2024-geographic], demonstrates the effectiveness of injecting geolinguistic signals directly into encoder architectures. Instead of relying on a single global classifier, which risks overgeneralization and false positives, GeoAdapt leverages region-conditioned plausible language sets [wang2025geolocationawarerobustspokenlanguage]. This strategy narrows the hypothesis space to dialects actually spoken in a given region, significantly reducing misclassification rates at scale. For example, in multilingual contexts such as Switzerland, studies on Swiss German reveal that while there is a robust and genuine dialectal signal detectable in embeddings, the proximity of closely related varieties leads to persistent confusion [vamvas-etal-2024-modular]. This highlights an important methodological pitfall: reporting only macro-averaged accuracy across languages obscures dialect-level failures and paints an overly optimistic picture of performance.
Dialect-sensitive modeling is equally crucial in settings involving code-switching. While LLMs now consistently outperform strong supervised baselines on benchmarked code-switching tasks [khanuja-etal-2020-gluecos], their performance deteriorates when confronted with more fluid, naturally occurring mixing patterns or when faced with domain shifts such as informal social media text. This brittleness underscores the need for stress-testing models under realistic multilingual conditions, rather than relying solely on curated datasets.
In the Arabic NLP community, the introduction of AraDiCE provides a concrete response to these challenges. Unlike many prior resources that implicitly privilege Modern Standard Arabic (MSA), AraDiCE explicitly pairs dialectal coverage with cultural knowledge [mousi-etal-2025-aradice]. This design avoids “standard-language centrism,” ensuring that dialects are treated as primary linguistic objects rather than deviations. Moreover, AraDiCE introduces quantifiable measures of coverage across dialect regions, allowing researchers to more transparently assess which varieties are well supported and which remain underrepresented [mousi-etal-2025-aradice].
@@
  • Audit and control for toponym shortcuts: models may learn to rely on explicit place names (“Cairo,” “Zurich”) rather than genuine linguistic cues. To mitigate this, it is preferable to use coarse regional metadata, and when necessary, remove place names altogether during training and evaluation [hofmann-etal-2024-geographic].
  • Incorporate code-switch stress testing by default: model evaluation should include tasks with rapid and fluid switching, reflecting authentic multilingual discourse rather than only balanced, neatly segmented datasets [khanuja-etal-2020-gluecos].
  • Stratify metrics within a language by variety or region: rather than reporting only cross-language means, which may obscure intra-language disparities, fine-grained stratification reveals performance trade-offs and highlights which dialects remain underserved [faisal-etal-24-dialectbenchy].
@@ 8.2.2 Internal mechanisms and cross-language generalization: a multilingual core and monolingual regions
Multilingual Transformers exhibit a sparse, modular organization: a compact multilingual core carries language-agnostic structure and meaning, while language-/family-specific regions capture orthography, morphology, and idiosyncratic constructions [foroutan-etal-2022-discovering, pires-etal-2019-multilingual, tang-etal-2024-language]. This motivates freeze-and-specialize: keep the core stable and attach small per-language adapters/LoRA or sparsely routed experts to absorb variation, reducing cross-language interference under fixed compute and mitigating the “curse of multilinguality” [pfeiffer-etal-2020-mad, shazeer2017moe, fedus2022switch, arivazhagan2019massively].
@@
  • Representations: cross-script pairs align in core layers; clustering follows syntax/semantics rather than script [pires-etal-2019-multilingual, chang2022geometry].
  • Causality: disabling locals hurts orthography/morphology with minor effect on cross-lingual semantics; core perturbations show the converse [pfeiffer-etal-2020-mad, foroutan-etal-2022-discovering].
  • Routing/sparsity: MoE gates differentiate languages; localized circuits emerge for language-specific phenomena [shazeer2017moe, fedus2022switch].
  • Dynamics: full finetuning forgets more; frozen cores + small locals yield a better stability–specialization trade-off [winata-etal-2023-overcoming].
@@
    Localize: fit small per-language & family modules; for MoE use top-$k$ routing and gate-entropy penalties; for adapters & LoRA enforce sparsity & low rank [shazeer2017moe, fedus2022switch, pfeiffer-etal-2020-mad].
@@
    Consolidate (optional): distill core and locals into a sparse expert layer to retain modularity at low latency [zhang-etal-2024-lightweight].
@@
Script & morphology engineering often dominates scale: (i) reversible, traceable normalization & transliteration with provenance [xhelili-etal-2024-breaking]; (ii) deliberate tokenization (byte & char & subword) by script and typology [yu2020pcgrad]; (iii) segmentation & reinflection as process-level checks so gains reflect structure, not orthography [shazeer2017moe, fedus2022switch, winata-etal-2023-overcoming].
@@ 9.1 Development of Specialized Language Models for Low-Resource Languages
    Few-shot and zero-shot learning provide avenues for adapting models to low-resource languages with minimal labeled data. Beyond common approaches like prompt-based learning, innovative frameworks such as Model-Agnostic Meta-Learning (MAML) [finn2017model] optimize models for task adaptability, enabling them to perform well with only a handful of examples. Dynamic task re-weighting can prioritize features unique to low-resource languages, ensuring better alignment with linguistic structures. Additionally, task-conditioning embeddings that incorporate linguistic typology can guide models to focus on crucial language-specific attributes. For zero-shot learning, retrieval-augmented reasoning can dynamically fetch relevant multilingual examples, improving the model’s context-aware adaptation to unseen languages.